OpenStudy (vvfvf):

Find the slope of the line passing through the points , −9−6 and , −45 .

OpenStudy (calculusxy):

Do you mean the points (-9, -6) and (-4, 5)?

OpenStudy (vvfvf):

yes

OpenStudy (calculusxy):

Use the formula to find the slope: \(\large \frac{y_2 -y_1}{x_2 - x_1} = m\)

OpenStudy (lukesolo_2187):

slope formula is y2−y1x2−x1=m

OpenStudy (calculusxy):

You can have -6 as \(y_2\) and -9 as \(x_2\). Then -4 can be \(x_1\) and 5 can be \(y_1\)

OpenStudy (lukesolo_2187):

5-(-6)/-4-(-9)

OpenStudy (lukesolo_2187):

11/5

OpenStudy (lukesolo_2187):

is the slope

OpenStudy (vvfvf):

i'd just got the same answe too thanks guys

OpenStudy (lukesolo_2187):

are u doing flvs

OpenStudy (lukesolo_2187):

can i get medal plz

OpenStudy (vvfvf):

do you mind helping me with another one plez?

OpenStudy (lukesolo_2187):

okay

OpenStudy (lukesolo_2187):

i am good with slopes

OpenStudy (lukesolo_2187):

sorry took a while to respond didnt notice

OpenStudy (vvfvf):

it's fine do you want me to type it here or do another post? i just need twomore then im good for now

OpenStudy (lukesolo_2187):

here is fine sorry again